Merchandise with a total price of $2,896 has discounts of 45% and is 15% and is sold with a 3% cash discount. How much is saved when discounts are taken.
Answer by addingup(3677)

Are the discounts applied to the total price or are they consecutive discounts?
I think they must be consecutive:
2896 = 100%
100-(100*0.45) = 55
55-(55*0.15)= 46.75
46.75-(46.75*0.03) = 45.35%
````````````````````````````````````````
2,896*0.4535 = 1,313.34 is the net price, and you save:
2,896-1,313.34 = 1,582.66
J